A player hits a mass kd 0.24kg moving northwards with a velocity of 15m/s thereby causing it to move with a velocity 20m/s southwards. The force of th blow acts on the ball for 0.025 . Calculate the average force exerted on the ball by the player.


1
Expert's answer
2021-11-07T19:27:55-0500

Determine the force:


"F=\frac{\Delta I}{t}=\frac{m\Delta v}{t}=\frac{m(v_f-v_i)}{t},\\\space\\
F=\frac{0.24(20-(-15))}{0.025}=336\text{ N}."